A solution of HCl is known to be 0.5622 N. When titrated against KOH, 45.22 ml of the acid neutralized 48.23 ml of the base. Calculate the N of the base.



 


1
Expert's answer
2022-05-30T17:02:02-0400

CN — concentration of normal;

V — volume of solution (litr);

E — equivalent amount of substance;

CN = E / V

1.

For HCl:

E = CN * V = 0,5622 * 0,04522 = 0,0254

substances react in equal equivalent amounts. therefore the KOH equivalent amount is also 0.0254

2.

For KOH:

CN = E / V = 0,0254 / 0,04823 = 0,527 N


ANSWER: CN (KOH) = 0,527 N